OpenStudy (anonymous):

When x is an odd number, the function f(x) = 2(x + 2) gives an even number. Identify the set of even numbers that correspond to this set of odd numbers that satisfy the function. {41, 57, 77, 83, 95} The set of even numbers that correspond to the given set of odd numbers that satisfy the function is

OpenStudy (anonymous):

hi.....just put these value instead of x...... f(41)=2(41+2)=2(43)=86 ....